OpenStudy (anonymous):

Please help. studying for my final and i don't really get functions… Determine which of the following are functions. \[A. \left| x \right|+2=y-1\] \[B. y ^{2}-2x=3\] \[c. 3x-y \le2\] \[D. 5x ^{2}-2x=y+1\]

OpenStudy (anonymous):

in these examples, do you want y as a function of x?

OpenStudy (anonymous):

well the question says determine which ones are functions so I'm not really sure.

OpenStudy (anonymous):

isn't there a way to look at the equation and know which one is a function?

OpenStudy (anonymous):

yes... solve for y in each... then determine that if there is more than 1 y value for a particular x value in the equation, then it is NOT a function...

OpenStudy (anonymous):

it looks like there is more than 1 answer according to the choices/examples you provided...

OpenStudy (anonymous):

let's try "A"... solve for y....

OpenStudy (anonymous):

ok so ill try to solve it and ill get back to you. and let you know what i did

OpenStudy (anonymous):

ok... tag me when u need help....

OpenStudy (anonymous):

ok thank you!

OpenStudy (anonymous):

@dpaInc already need help but what i did was to get y by itself and got is \[\left| x \right|+3=y\] but unsure what my next step is.

OpenStudy (anonymous):

ok.. that's fine...

OpenStudy (anonymous):

notice here, that for ANY x value, you will get a unique y value.... so is this a function?

OpenStudy (anonymous):

do you know about the vertical line test?

OpenStudy (anonymous):

yes it is a function right? since there could be multiple x's but not y

OpenStudy (anonymous):

correct... this is a function....

OpenStudy (anonymous):

do you know of the vertical line test?

OpenStudy (anonymous):

the points can't hit the line more than once right? and if it does hit it more than once it is not a function

OpenStudy (anonymous):

yes... and that's the way i do these problems... with a picture in my head... so for example... in A, i know the graph looks like this. and if i draw a vertical line anywhere, the vertical line will hit the graph only at 1 point...

OpenStudy (anonymous):

so for the last problem since it was an absolute value is that what made it a function?

OpenStudy (anonymous):

oh ok so i can just type it into my calculator and look at each problem?

OpenStudy (anonymous):

no... even without the absolute value, this would still be a function... BUT, if the absolute value was with the y variable, then it would not be a function....

OpenStudy (anonymous):

so anytime the y is an absolute value it is automatically not a function?

OpenStudy (anonymous):

yes...

OpenStudy (anonymous):

ok and how about for exponents? are there any rules for those

OpenStudy (anonymous):

let's do an easy example... which of these is a function: \(\huge y=x^2 \) \(\huge x=y^2 \)

OpenStudy (anonymous):

y=x^2 right?

OpenStudy (anonymous):

correct... why is that second one NOT a function?

OpenStudy (anonymous):

because X is one number and y can equal multiple numbers

OpenStudy (anonymous):

right...:)

OpenStudy (anonymous):

so do you think your second example in your question is a function?

OpenStudy (anonymous):

ok cool! so for my problems b is not a function and c is a function and still working on d

OpenStudy (anonymous):

correct...

OpenStudy (anonymous):

and d is a function as well right?

OpenStudy (anonymous):

here's the thing with "c".... it is not an equation but it is an inequality... that means you'll need to shade above or below some curve (or line, in this case) to show the solution set.... will the vertical line hit more than 1 y-value in the solution set if you do the vertical line test?

OpenStudy (anonymous):

you're correct with "d"....

OpenStudy (anonymous):

for c either side i shade it will hit the y value more than once meaning it will not be a function, correct?

OpenStudy (anonymous):

yes...:) you are correct....

OpenStudy (anonymous):

so any time i get a inequality it will not be a function?

OpenStudy (anonymous):

yes.... but there are exceptions... may i ask, what class is this? algebra 1?

OpenStudy (anonymous):

its intermediate algebra for the junior college

OpenStudy (anonymous):

ok... that's fine...

OpenStudy (anonymous):

you'll probably get to the part where I said "there are exceptions" later on in this course...:)

OpenStudy (anonymous):

can we recap? which are functions, which are not?

OpenStudy (anonymous):

i am at the end of the course so maybe the next class. the problems i posted were off my study guide for the final.

OpenStudy (anonymous):

so A is a function, B and C are not, and D is a function!

OpenStudy (anonymous):

yes... good luck on your final...:) the things we did here you should definitely write down....

OpenStudy (anonymous):

looks good...:)

OpenStudy (anonymous):

I am! thank you so much for your help! i understand functions better now!

OpenStudy (anonymous):

yw...:)